You have one type of nut that sells for $2.70/lb and another type of nut that sells for $6.40/lb. You would like to have 22.2 lbs of a nut mixture that sells for $5.40/lb. How much of each nut will you need to obtain the desired mixture?
You will need
lbs of the cheaper nut
and
lbs of the expensive nut.

Answer by Alan3354(69443)

Do it like this one. Only the numbers are different.
-----------
Amanda wants to make 8 gal. of a 20% saline solution by mixing together a 56% saline solution and a 8% saline solution. How much of each solution must she use?
================
e = amount of 8%
f = amount of 56%
---
e+ f = 8 (total solution)
8e + 56f = 20*8 (total saline)
---
e+ f = 8
e + 7f = 20
------------ Subtract
-6f = -12
f = 2
e = 6